OpenStudy (anonymous):

nick1234567 Area of one leaf of the rose r=3sin(4theta)

OpenStudy (michele_laino):

I'm very sorry, it is 2:38 a.m. (Italy time zone) from me so I have to go to sleep

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

Need answer in next 3 mins please

OpenStudy (anonymous):

Ur the best!

pooja195 (pooja195):

ok any ideas on where to start?

OpenStudy (anonymous):

Nope no clue and it's my last prob due in 3 mins

OpenStudy (anonymous):

Can u please save me

OpenStudy (anonymous):

So? Area is?

OpenStudy (anonymous):

Only have 1min left thank u again

OpenStudy (anonymous):

Ideas???

OpenStudy (anonymous):

Needs suggestion in the ext 39 sec

pooja195 (pooja195):

Things like this cant be done in that much time. You need to make sure you have time for it

pooja195 (pooja195):

You will need to use \[\huge~\rm~\int\limits_{}^{}1/2r^2d \emptyset\]

OpenStudy (irishboy123):

if this is an emergency, this is a brief attempt: \(r=3sin(4 \theta)\) \(r = 0, \ 4\theta = 0, \ \pi \implies \theta = 0, \pi/4, ....\) \(A = \frac{1}{2} \int_{0}^{\pi / 4} 9 sin^2 (4 \theta) \ d \theta \) \( = \frac{9}{4} \ \int_{0}^{\pi / 4} 1 - cos \ 8 \theta \ d \theta\) \( = \frac{9}{4} \ [ \theta - \frac{1}{8}sin \ 8 \theta \ ]_{0}^{\pi / 4}\) \( = \frac{9}{4} \ [ \frac{\pi}{4} ] = \frac{9 \pi}{16}\)

pooja195 (pooja195):

^or that works too ;p

pooja195 (pooja195):

You will need to use \[\huge~\rm~\int\limits_{}^{}1/2r^2d \theta\] \[\huge~\rm~\int\limits\limits_{\alpha }^{\beta }1/2r^2d \theta \] \[\huge~\rm~\int\limits\limits\limits_{0 }^{\ \pi/4 }1/2(3\sin(4\theta))^2d \theta =\frac{ 9 }{ 16 }\pi \]

OpenStudy (jack1):

lol... why is this q orange?